Solve the exponential equation.
1/16= 64^(4x–3)

Answer by lwsshak3(11628)

Solve the exponential equation.
1/16= 64^(4x–3)
1/4^2=4^3^(4x-3)
4^-2)=4^(12x-9)
12x-9=-2
12x=7
x=7/12